OpenStudy (anonymous):

if y5 = 25 and x2 = 5 , what is the value of y10x-2?

OpenStudy (kira_yamato):

\[y^5=25,x^2=5\] \[y=\sqrt[5]{25}, x=\sqrt{5}\] \[y=5^{2/5}, x = 5^{1/2}\] I'll leave you with this...

OpenStudy (anonymous):

how i can Solve that ? :( by these